hartnn (hartnn):

Is 0 a polynomial ? what is the degree ? I said YES, '0' is a polynomial and f(x)=0 has degree \(-\infty\)

hartnn (hartnn):

am i correct?

OpenStudy (lgbasallote):

it has a degree of 0

hartnn (hartnn):

sure?? just because its constant ?

OpenStudy (lgbasallote):

only \(e^{-\infty}\) equals 0..the rest of the variables don't

OpenStudy (lgbasallote):

well the degree is also infinite

OpenStudy (lgbasallote):

0x^2 = 0

hartnn (hartnn):

so whats the final degree? i have read it somewhere that its \(-\infty\) or -1

mathslover (mathslover):

undefined ..?

mathslover (mathslover):

http://mathworld.wolfram.com/ZeroPolynomial.html

OpenStudy (lgbasallote):

personally i think it's indeterminate :|

hartnn (hartnn):

yeah, i have read there only, wolfram

mathslover (mathslover):

some crazy authors do say that given in the link .. but it is basically undefined

mathslover (mathslover):

The degree of the zero polynomial is either left undefined, or is defined to be negative (usually −1 or −∞). --> wikipedia http://en.wikipedia.org/wiki/Degree_of_a_polynomial

mathslover (mathslover):

http://planetmath.org/ZeroPolynomial2.html

hartnn (hartnn):

hmm...thanks for the replies,though.

hartnn (hartnn):

but everyone is sure about 0 being a polynomial, right ?

OpenStudy (anonymous):

Sorry, but, you should not define \[ \deg(0)=0 \]Typically, it's a good idea to define it as: \[ \deg(0)=-\infty \]But, it is mainly used for well-ordering principles in rings \(R\), such that it is \(R[x]\). It can also be left undefined, but either definition causes problems at some point or another in a proof.

OpenStudy (anonymous):

And I say "well-ordering principles" with a grain of salt. I mean it to be something similar to the applied well-ordering of \(\mathbb{Z}\), or absolute ordering of \(\mathbb{Q}\) or \(\mathbb{R}\).

OpenStudy (anonymous):

Last random, separated post... hopefully. So, we take, for example: \[ P(x), Q(x)\in R[x] \]Where \(R[x]\) is a polynomial ring, then, we wish to maintain the following desirable properties (required for any absolute measure): \[ \deg(P(x)+Q(x))\leq\max(\deg(P(x)), \deg(Q(x)))\\ \deg(P(x)Q(x))\leq \deg(P(x))+\deg(Q(x)) \]That's why it leads us to the conclusion of \(\deg(0)=-\infty\) (where \(0\in R\) is the additive identity). The point is, though, it *could* screw you over, either way, in a proof (as it has me, more than once in number theory), unless you define that specific case as something different, or treat it as a separate case.

hartnn (hartnn):

really appreciate that ^^ thanks :)

OpenStudy (anonymous):

Sure thing... great question, by the way.